# AI4Science

### v1

@@ 2 Stellarator Design: Preliminaries, Problem Set-up, and Related Work
Stellarator design is focused, in part, on manipulating properties of magnetic fields with nested surfaces. Points on a magnetic surface can be described with two angles: the poloidal angle, $\theta$, and toroidal angle, $\varphi$, see Figure˜1B; the toroidal angle varies the long way around the torus while the poloidal angle varies the short way around the torus. A magnetic surface has three basic attributes: the number of field periods $n_{\text{fp}}$, the aspect ratio AA, and the rotational transform, $\iota$. The number of field periods, $n_{\text{fp}}$, refers to a discrete rotational symmetry, i.e., rotating the device by $2\pi/n_{\text{fp}}$ about the $Z$-axis results in the same device. A device with $n_{\text{fp}}$-field periods can be sliced (like a pizza) into $n_{\text{fp}}$ identical pieces; for example, Fig.˜3 (top row) shows devices with $n_{\text{fp}}=2,3,4,5,6,7,8$, from left to right. The aspect ratio, $A=R/r$, is a geometric quantity measuring the ratio of the major to minor radius of the device, see Figure˜1B, and reflects how thin the enclosed plasma is. Low aspect ratio devices are often preferred, since they correspond to a high volume of plasma confined, relative to the footprint. A third property which characterizes magnetic surfaces is the rotational transform $\iota$. The rotational transform is the average number of poloidal rotations a field line makes for every toroidal turn around the torus. Figure˜1B illustrates a hypothetical red field line making, roughly, one poloidal turn as it transits one field period, leading to $\iota\approx n_{\text{fp}}=2$. When given multiple nested magnetic surfaces with different values of $\iota$, the average value, $\overline{\iota}$, is considered. When designing stellarators, the shape of $S$ is varied to achieve desirable values of $\overline{\iota},A$.
@@
In this work, we focus on solutions to the stage-I problem. The decision variables, $\mathbf{x}$, parameterize the shape of the plasma boundary, $S$. Two common constraints, which we consider in this work, constrain the aspect ratio and $\overline{\iota}$ of the device to target values $A^{*},\overline{\iota}^{*}$: $c_{A}(\mathbf{x})=\frac{A(\mathbf{x})-A^{*}}{A^{*}}$ and $c_{\iota}(\mathbf{x})=\frac{\overline{\iota}(\mathbf{x})-\overline{\iota}^{*}}{\overline{\iota}^{*}}$ landreman_magnetic_2022; baillod_stellarator_2022; baillod_integrating_2025; jorge_single-stage_2023; landreman_mapping_2022. The objective function used in this work, $J=J_{QS}$ giuliani_direct_2022, measures the deviation of the magnetic field from the nearest quasisymmetric magnetic field, and is a proxy measure for particle confinement (see Appendix˜A). Quasisymmetry is a symmetry in the strength of the magnetic field, $\|\mathbf{B}\|$, that ensures that particle trajectories have no radial (outward) drift, thus ensuring good particle confinement helander_theory_2014; landreman_magnetic_2022. Quasisymmetric magnetic fields can be identified by plotting the contours of the field strength in a special coordinate system, Boozer coordinates – the field strength contours of quasisymmetric magnetic fields only depend on a linear combination of the Boozer angles, $\|\mathbf{B}\|(\theta,\varphi)=\|\mathbf{B}\|(\theta-Nn_{\text{fp}}\varphi)$, thus appearing straight in Boozer coordinates. For example, in Fig.˜1(C) the fieldlines are approximately parallel to the $x$-axis (aside from the small curve near the center). The type of quasisymmetry is parameterized by the normalized helicity, NN: $N=0$ denotes quasiaxisymmetry (QA) in which $\|\mathbf{B}\|$ appears flat when plotted in Boozer coordinates e.g. Fig.˜1(C); $N=1$ denotes quasi-helical symmetry (QH) in which the contours of $\|\mathbf{B}\|$ form diagonal lines when plotted in Boozer coordinates i.e. Fig.˜3(second row, column 3). Both types of quasisymmetry are desirable, however, there is an interplay between $n_{\text{fp}}$ and NN that makes it difficult to find high $n_{\text{fp}}$, QA devices, and low $n_{\text{fp}}$, QH devices. The metrics $c_{\iota},c_{A},J_{QS}$ can all be computed from the solution of Equation˜1, see Appendix A for details.
@@ 3 Case Study: Designing Stellarators With Diffusion Models
Training data We use data from the QUAsisymmetric Stellaratator Repository (QUASR). With over 300,000 stellarators, the QUASR repository is a large database of optimized stellarators including both coils and magnetic surfacs. QUASR was generated using Bayesian global optimization over the space of vacuum (current-free) stellarators, in search of quasisymmetric stellarators with a range of $n_{\text{fp}}$, aspect ratio AA, and rotational transform $\iota$ giuliani_comprehensive_2025; giuliani_direct_2024.
@@ Appendix A Description and Evaluation of Stellarator Designs
In this section, we discuss the representation of surfaces and evaluation of $J_{QS},c_{A},c_{\iota}$. Surfaces can be constructed with SIMSOPT [landreman_simsopt_2021] and evaluated with VMEC. Some stellarators in the training data, or generated by the diffusion model, have highly shaped boundaries. VMEC struggles to evaluate these, failing to return a solution for approximately $39\%$ of stellarators from the training data, and $49\%$ of stellarators generated in-sample by the diffusion model. In reporting results, we ignore these designs.
@@
The metrics $J_{QS},c_{\iota},c_{A}$ are evaluated with the following protocol. Given $n_{\text{fp}}$ and $\mathbf{x}$ a surface is object is constructed in SIMSOPT. The surface is evaluated with VMEC, from which $\overline{\iota}$ and AA can be computed. Boozer coordinates are constructed using the BoozXform in SIMSOPT. $B$ and the jacobian $\sqrt{g}$ are constructed in Boozer coordinates, from which $J_{QS}$ can be evaluated.
@@ Appendix B Diffusion Model Details
The model was trained with ${n_{\mathbf{y}}}=4$ features, $\mathbf{y}=(\overline{\iota},A,n_{\text{fp}},N)$. The diffusion model used $200$ timesteps, with a linear schedule. A standard feed-forward neural network architecture was used to describe $f_{\theta}(\mathbf{x}_{t},t,\mathbf{y})$ with 44 hidden layers, each with width $2048$, and Gaussian error linear unit (GELU) activiation functions. A sinusiodal network head was used to embed $\mathbf{x},t,\mathbf{y}$ before entering the standard feed-forward layers. The network head maps $\mathbf{x}$ to dimension $64$, $t$ to dimension $128$, and $\mathbf{y}$ to dimension $128$. The Adam optimizer [kingma_adam_2017] was used with a training batch size of $4096$, an evaluation batch size of $128$, for $250$ epochs, with a learning rate of $0.0005$.

### v2

@@ 2 Stellarator Design: Preliminaries, Problem Set-up, and Related Work
Stellarator design is focused, in part, on manipulating properties of magnetic fields with nested surfaces. Points on a magnetic surface can be described with two angles: the poloidal angle, $\theta$, and toroidal angle, $\varphi$, see Figure˜1B; the toroidal angle varies the long way around the torus while the poloidal angle varies the short way around the torus. A magnetic surface has three basic attributes: the number of field periods $n_{\text{fp}}$, the aspect ratio $A$, and the rotational transform, $\iota$. The number of field periods, $n_{\text{fp}}$, refers to a discrete rotational symmetry, i.e., rotating the device by $2\pi/n_{\text{fp}}$ about the $Z$-axis results in the same device. A device with $n_{\text{fp}}$-field periods can be sliced (like a pizza) into $n_{\text{fp}}$ identical pieces; for example, Fig.˜3 (top row) shows devices with $n_{\text{fp}}=2,3,4,5,6,7,8$, from left to right. The aspect ratio, $A=R/r$, is a geometric quantity measuring the ratio of the major to minor radius of the device, see Figure˜1B, and reflects how thin the enclosed plasma is. Low aspect ratio devices are often preferred, since they correspond to a high volume of plasma confined, relative to the footprint. A third property which characterizes magnetic surfaces is the rotational transform $\iota$. The rotational transform is the average number of poloidal rotations a field line makes for every toroidal turn around the torus. Figure˜1B illustrates a hypothetical red field line making, roughly, one poloidal turn as it transits one field period, leading to $\iota\approx n_{\text{fp}}=2$. When given multiple nested magnetic surfaces with different values of $\iota$, the average value, $\overline{\iota}$, is considered. When designing stellarators, the shape of $S$ is varied to achieve desirable values of $\overline{\iota},A$.
@@
In this work, we focus on solutions to the stage-I problem. The decision variables, $\mathbf{x}$, parameterize the shape of the plasma boundary, $S$. Two common constraints, which we consider in this work, constrain the aspect ratio and $\overline{\iota}$ of the device to target values $A^{*},\overline{\iota}^{*}$: $c_{A}(\mathbf{x})=\frac{A(\mathbf{x})-A^{*}}{A^{*}}$ and $c_{\iota}(\mathbf{x})=\frac{\overline{\iota}(\mathbf{x})-\overline{\iota}^{*}}{\overline{\iota}^{*}}$ landreman_magnetic_2022; baillod_stellarator_2022; baillod_integrating_2025; jorge_single-stage_2023; landreman_mapping_2022. The objective function used in this work, $J=J_{QS}$ giuliani_direct_2022, measures the deviation of the magnetic field from the nearest quasisymmetric magnetic field, and is a proxy measure for particle confinement (see Appendix˜A). Quasisymmetry is a symmetry in the strength of the magnetic field, $\|\mathbf{B}\|$, that ensures that particle trajectories have no radial (outward) drift, thus ensuring good particle confinement helander_theory_2014; landreman_magnetic_2022. Quasisymmetric magnetic fields can be identified by plotting the contours of the field strength in a special coordinate system, Boozer coordinates – the field strength contours of quasisymmetric magnetic fields only depend on a linear combination of the Boozer angles, $\|\mathbf{B}\|(\theta,\varphi)=\|\mathbf{B}\|(\theta-Nn_{\text{fp}}\varphi)$, thus appearing straight in Boozer coordinates. For example, in Fig.˜1(C) the fieldlines are approximately parallel to the $x$-axis (aside from the small curve near the center). The type of quasisymmetry is parameterized by the normalized helicity, $N$: $N=0$ denotes quasiaxisymmetry (QA) in which $\|\mathbf{B}\|$ appears flat when plotted in Boozer coordinates e.g. Fig.˜1(C); $N=1$ denotes quasi-helical symmetry (QH) in which the contours of $\|\mathbf{B}\|$ form diagonal lines when plotted in Boozer coordinates i.e. Fig.˜3(second row, column 3). Both types of quasisymmetry are desirable, however, there is an interplay between $n_{\text{fp}}$ and $N$ that makes it difficult to find high $n_{\text{fp}}$, QA devices, and low $n_{\text{fp}}$, QH devices. The metrics $c_{\iota},c_{A},J_{QS}$ can all be computed from the solution of Equation˜1, see Appendix A for details.
@@ 3 Case Study: Designing Stellarators With Diffusion Models
Training data We use data from the QUAsisymmetric Stellaratator Repository (QUASR). With over 300,000 stellarators, the QUASR repository is a large database of optimized stellarators including both coils and magnetic surfacs. QUASR was generated using Bayesian global optimization over the space of vacuum (current-free) stellarators, in search of quasisymmetric stellarators with a range of $n_{\text{fp}}$, aspect ratio $A$, and rotational transform $\iota$ giuliani_comprehensive_2025; giuliani_direct_2024.
@@ Appendix A Description and Evaluation of Stellarator Designs
In this section, we discuss the representation of surfaces and evaluation of $J_{QS},c_{A},c_{\iota}$. Surfaces can be constructed with SIMSOPT [landreman_simsopt_2021] and evaluated with VMEC. Some stellarators in the training data, or generated by the diffusion model, have highly shaped boundaries. VMEC struggles to evaluate these, failing to return a solution for approximately $41\%$ of stellarators from the training data, and $57\%$ of stellarators generated in-sample by the diffusion model. In reporting results, we ignore these designs.
@@
The metrics $J_{QS},c_{\iota},c_{A}$ are evaluated with the following protocol. Given $n_{\text{fp}}$ and $\mathbf{x}$ a surface is object is constructed in SIMSOPT. The surface is evaluated with VMEC, from which $\overline{\iota}$ and $A$ can be computed. Boozer coordinates are constructed using the BoozXform in SIMSOPT. $B$ and the jacobian $\sqrt{g}$ are constructed in Boozer coordinates, from which $J_{QS}$ can be evaluated.
@@ Appendix B Diffusion Model Details
The model was trained with ${n_{\mathbf{y}}}=4$ features, $\mathbf{y}=(\overline{\iota},A,n_{\text{fp}},N)$. The diffusion model used $200$ timesteps, with a linear schedule. A standard feed-forward neural network architecture was used to describe $f_{\theta}(\mathbf{x}_{t},t,\mathbf{y})$ with $4$ hidden layers, each with width $2048$, and Gaussian error linear unit (GELU) activiation functions. A sinusiodal network head was used to embed $\mathbf{x},t,\mathbf{y}$ before entering the standard feed-forward layers. The network head maps $\mathbf{x}$ to dimension $64$, $t$ to dimension $128$, and $\mathbf{y}$ to dimension $128$. The Adam optimizer [kingma_adam_2017] was used with a training batch size of $4096$, an evaluation batch size of $128$, for $250$ epochs, with a learning rate of $0.0005$.